### v1

# Dirac Neutrinos and Dark Matter within a Minimal Discrete Symmetry Model

Dirac Neutrinos and Dark Matter within a Minimal
Discrete Symmetry Model
[…]
aSchool of Physical Science and Technology, Xinjiang University, Urumqi 830017, China
We present a model that extends the standard model by incorporating the simplest discrete symmetry groups, $Z_{2}$ and $Z_{3}$. This model introduces vector-like leptons and a real scalar singlet. Based on this framework, we generate Dirac neutrino masses and explain the neutrino normal mass ordering. The model also aligns well with current oscillation data regarding theoretical values of the leptonic mixing angles and the Dirac CP-violating phase. Furthermore, it predicts that the atmospheric mixing angle falls in the higher octant and proposes a viable dark matter candidate. We discuss the phenomenological aspects and future testability of the model.
1 Introduction
Neutrino mass and dark matter represent significant drawbacks of the Standard Model (SM). Addressing either of these requires an extension of the SM through the construction of a new physics model. The experimental discovery of neutrino oscillations serves as a clear indicator of non-zero neutrino mass and mixing [1, 2]. However, several aspects remain unresolved, including the nature of the neutrino (Dirac or Majorana), mass ordering, precise values of some mixing angles, and the leptonic CP-violating phase. Similarly, while there is overwhelming observational evidence for the existence of dark matter (DM) [3, 4, 5], questions about its particle nature and production mechanisms, such as thermal freeze-out [6, 7, 8, 9] or freeze-in [10, 11, 12], persist. Neutrinos and DM share characteristics as weakly interacting, invisible, and elusive particles, making it compelling to unify them within a single theoretical framework and explore potential dual-purpose solutions.
The distinction between Dirac and Majorana neutrinos leads to two different mass generation mechanisms. Neutrinos are often expected to be Majorana fermions, primarily because they are color and charge neutral, and the model-independent effective field theory approach—namely, the seesaw mechanism—most naturally and elegantly generates Majorana masses. However, the absence of experimental confirmation from neutrinoless double beta decay experiments, despite considerable efforts, suggests it is premature to dismiss the possibility of Dirac neutrinos. Recent studies have shown that the seesaw mechanism can be adapted to generate Dirac neutrino masses at the tree level, referred to as the Dirac seesaw [13, 14, 15, 16, 17, 18, 19, 20]. An intriguing feature of this mechanism is that it also simultaneously addresses other beyond-SM issues, such as DM [21, 22, 23, 24, 25, 26, 20, 27, 28], the strong CP problem [29, 30, 31, 32, 33, 34], and leptogenesis [35, 36, 37, 38, 39, 40]. The recent surge of interest in the dynamical generation of Dirac neutrino mass, alongside simultaneous solutions for other outstanding challenges to the SM, highlights critical features testable by experiments.
Given their feebly interacting properties, the underlying theories of DM and neutrino mass generation may be interconnected. In certain models, neutrinos play a role analogous to DM, although they are generally considered distinct. Notably, some models posit DM as essential for mediating interactions necessary for neutrino mass generation, such as in scotogenic [41, 42], dark linear [43, 44], and inverse seesaw mechanisms [45]. Moreover, in Dirac seesaw mechanisms, the symmetry envisioned to preserve lepton number and prohibit Majorana mass terms also stabilizes the DM [21, 22, 23, 24, 25, 26, 20, 27, 28]. This pursuit of a novel model wherein Dirac neutrino mass is generated by the seesaw mechanism with heavy fermionic DM in the internal line is of great scientific importance.
In pursuit of simplicity and minimalism, this paper aims to construct an unexplored Dirac seesaw model, generate neutrino masses and mixing, and account for the dark matter relic abundance. The structure of this paper is organized as follows. Section 2 details the model construction, focusing on symmetries, field content, the Yukawa interaction Lagrangian, which generates neutrino and charged lepton masses, and the scalar potential, resulting in scalar masses and their interactions. Phenomenological consequences are discussed in section 3, where we show that the model explains neutrino normal mass ordering, oscillation parameters, and DM relic abundance, and also provides predictions. Section 4 presents concluding remarks.
2 The model
It is well-known that the SM, being a low-energy effective theory, is incomplete and must be augmented at higher energies. In what follows, we try to construct such a UV theory based on guiding principles of symmetry. In adherence to the symmetry principle and the seesaw idea, our main goal in this section is to forbid the usual Majorana mass term for the right-handed neutrino, while achieving the Dirac mass term $m_{D}\overline{\nu_{L}}\nu_{R}$ after integrating out the heavy states. To this end, we introduce a model extending the SM by incorporating $Z_{2}\times Z_{3}$ symmetries and by adding three generations of right-handed neutrinos $\nu_{R}$ and vector-like lepton doublets $\psi_{L,R}=(\psi^{0}_{L,R},\psi^{-}_{L,R})^{T}$. Additionally, the scalar sector includes a real singlet $\sigma$, besides the Higgs doublet $H$. The electroweak gauge quantum numbers and the transformation properties under the discrete symmetries are specified in Table 1. As indicated, all leptons transform similarly under the $Z_{3}$ symmetry, while the scalars are invariant. For the $Z_{2}$ symmetry, the SM fields are invariant, whereas the newly introduced fields change sign.
Fields $L_{i}$ $e_{i,R}$ $\nu_{i,R}$ $\psi_{i,L}$ $\psi_{i,R}$ $H$ $\sigma$
$SU(2)_{\mathrm{L}}$ $2$ $1$ $1$ $2$ $2$ $2$ $1$
$U(1)_{\mathrm{Y}}$ $-\frac{1}{2}$ $-1$ 00 $-\frac{1}{2}$ $-\frac{1}{2}$ $\frac{1}{2}$ 00
$Z_{3}$ $\omega$ $\omega$ $\omega$ $\omega$ $\omega$ $1$ $1$
$Z_{2}$ $+1$ $+1$ $-1$ $-1$ $-1$ $1$ $-1$
Table 1: The electroweak $SU(2)_{\mathrm{L}}\times U(1)_{\mathrm{Y}}$ gauge quantum numbers and $Z_{2}$, $Z_{3}$ transformation properties of the left-handed lepton doublets $L_{i}$ (with $i=1,2,3$), the right-handed neutrinos $\nu_{i,R}$, the vector-like leptons $\psi_{i,L}$ and $\psi_{i,R}$, the Higgs doublet $H$ and the scalar singlet $\sigma$. Here, $\omega$ is a cubic root of unity, e.g. $\omega=e^{2\pi i/3}$.
The Yukawa interactions between the scalars and leptons are expressed in the following Lagrangian:
$-\mathcal{L}_{Y}=Y_{E,ij}\bar{L}_{i}He_{j,R}+Y_{\psi,ij}\bar{L}_{i}\psi_{j,R}\sigma+Y_{\nu,ij}\overline{\psi_{i,L}}\tilde{H}\nu_{j,R}+\mathrm{h.c.},$ (2.1)
where $\tilde{H}=i\sigma^{2}H^{*}$. Although the Yukawa interaction $\bar{L}\tilde{H}\nu_{R}$, which is needed for the neutrino Dirac mass term, is allowed by $Z_{3}$ symmetry, it is forbidden by the $Z_{2}$. This mass term is generated when the latter is spontaneously broken by vacuum expectation value (vev) of $\sigma$ and after the heavy vector-like leptons are integrate out. Consequently, the small neutrino mass is explained. The $Z_{3}$ symmetry disallows the Majorana mass term for the right-handed neutrino, even after the scalars gain vevs, thus preserving the Dirac nature of the neutrino. The mass term for the vector-like leptons remains invariant under all symmetries and reads $M_{ij}\overline{\psi_{i,L}}\psi_{j,R}$. This mass term is expected to be large, which suppresses the neutrino mass via the seesaw mechanism, depicted in Figure 1.
Figure 1: The tree level Feynmann diagram for neutrino mass generation.
The scalar potential, invariant under both SM and the extended symmetries $Z_{2}\times Z_{3}$, is given by
$V(H,\sigma)=-\mu_{H}^{2}H^{\dagger}H-\frac{1}{2}\mu_{\sigma}^{2}\sigma^{2}+\frac{\lambda_{H}}{2}(H^{\dagger}H)^{2}+\frac{\lambda}{2}(H^{\dagger}H)\sigma^{2}+\frac{\lambda_{\sigma}}{4}\sigma^{4}.$ (2.2)
All couplings in this potential can be made real. The potential being bounded from below requires that the quartic couplings satisfy $\lambda_{H}\geq 0,\lambda_{\sigma}\geq 0$ and $\lambda+\sqrt{2\lambda_{H}\lambda_{\sigma}}\geq 0$. Expanding the scalar fields around their vevs, $H=(G^{+},(v+H^{0}+iG^{0})/\sqrt{2})^{T}$, $\sigma=(u+\sigma^{\prime})/\sqrt{2}$, we obtain stationary conditions for the potential:
$\mu^{2}_{H}=$ $\frac{1}{4}\left(2\lambda_{H}v^{2}+\lambda u^{2}\right),$ (2.3)
$\mu^{2}_{\sigma}=$ $\frac{1}{2}\left(\lambda v^{2}+\lambda_{\sigma}u^{2}\right).$
Given the expansions of the scalar expansions, we rewrite the potential in terms of the neutral and charged scalar components and subsequently extract the mass matrix for the two neutral scalars,
$M^{2}_{H^{0},\sigma^{\prime}}=\begin{pmatrix}\lambda_{H}v^{2}&\frac{1}{2}\lambda uv\\ \\[0.1pt] \frac{1}{2}\lambda uv&\frac{1}{2}\lambda_{\sigma}u^{2}\end{pmatrix}.$ (2.4)
This mass matrix is diagonalized by a $2\times 2$ orthogonal rotation matrix characterized by an angle $\theta$, which satisfies the relation $\tan 2\theta=\frac{2\lambda uv}{2\lambda_{H}v^{2}-\lambda_{\sigma}u^{2}}$. Diagonalization yields two physical states, $h$ and $\phi$, with the respective masses given by
$m^{2}_{h}=$ $\frac{1}{4}\left[2\lambda_{H}v^{2}+\lambda_{\sigma}u^{2}-\sqrt{\left(2\lambda_{H}v^{2}-\lambda_{\sigma}u^{2}\right)^{2}+(2\lambda uv)^{2}}\,\right],$ (2.5)
$m^{2}_{\phi}=$ $\frac{1}{4}\left[2\lambda_{H}v^{2}+\lambda_{\sigma}u^{2}+\sqrt{\left(2\lambda_{H}v^{2}-\lambda_{\sigma}u^{2}\right)^{2}+(2\lambda uv)^{2}}\,\right].$
The transformations between the original scalar fields and the mass eigenstates are governed by the following linear relationships:
$h=$ $\cos\theta H^{0}+\sin\theta\sigma^{\prime},$ (2.6)
ϕ=italic-$\phi=$ $\cos\theta\sigma^{\prime}-\sin\theta H^{0}.$
We identify $h$ as the SM-like Higgs boson with a mass of $m_{h}=125$ GeV, and $\phi$ as another neutral scalar, which is heavier than the Higgs.
The expression for the mixing angle $\theta$ and the scalar potential parameter $\lambda$, as shown in eq. (2.2), are crucial in determining the degree of mixing between these two scalars. Notably, in the absence of $\lambda$, there is no mixing ($\theta=0$), leading the scalars to decouple, as anticipated. On the other hand, the same parameter determines whether or not these scalar masses degenerate. Specifically, if $\lambda=0$, there exists a parameter space where the relation for two quartic couplings $\lambda_{\sigma}=2(v^{2}/u^{2})\lambda_{H}$ holds, resulting in equal mass squares, $m^{2}_{h}=m^{2}_{\phi}=\lambda_{H}v^{2}$. Conversely, when $\lambda$ is nonzero, it is not possible to achieve mass equality. If $\lambda\neq 0$, the minimum mass-squared difference, $m^{2}_{\phi}-m^{2}_{h}=\lambda uv$, occurs under the condition that the aforementioned relationship between $\lambda_{\sigma}$ and $\lambda_{H}$ is maintained. However, mass degeneracy is prohibited in this scenario. Additionally, the vacuum expectation values $u$ (also $v$) influence the masses of these scalars. For instance, if $v\gg u=0$, $m_{h}$ vanishes, and resulting in $m^{2}_{\phi}=\lambda_{H}v^{2}>m^{2}_{h}$. Conversely, when $u\gg v$, the mass squared of $h$ becomes $m^{2}_{h}\simeq(\lambda_{H}-\frac{\lambda^{2}}{2\lambda_{\sigma}})v^{2}+\mathcal{O}(v^{2}/u^{2})$, which is significantly smaller than $m^{2}_{\phi}\simeq\frac{1}{2}(\lambda_{\sigma}u^{2}+\frac{\lambda^{2}}{\lambda_{\sigma}}v^{2})+\mathcal{O}(v^{2}/u^{2})$, assuming all quartic couplings are of similar magnitude. Therefore, one can qualitatively expect the mass of $\phi$ to be substantially larger than that of the Higgs boson without detailed analysis of these couplings. Given the definition of the mass basis, the alternative assignment of the heavier scalar as the SM Higgs and the lighter one as the new neutral scalar is also valid. A similar discussion can be applied to this configuration. Additionally, other would-be Nambu-Goldstone bosons, i.e., charged and pseudo-scalar components of the Higgs doublet, remain massless and become the longitudinal components of the $W$ and $Z$ bosons.
Regarding the fermionic sector of the model, masses are generated after the spontaneous breaking of electroweak and extended symmetries. The charged lepton mass matrix, in the bases of $\left(\overline{e_{i,L}},\overline{\psi^{-}_{i,L}}\right)$ and $(e_{j,R},\psi^{-}_{j,R})^{T}$, is given by
$M_{E}=\begin{pmatrix}m_{E}&m_{\psi}\\ 0&M\end{pmatrix},$ (2.7)
where each entry consists of $3\times 3$ block matrices $m_{E}\equiv\frac{1}{\sqrt{2}}Y_{E}v$, $m_{\psi}\equiv\frac{1}{\sqrt{2}}Y_{\psi}u$ and $M$. There is a mixing between the left-handed charged lepton and the right-handed component of the vector-like lepton, while no mixing occurs between the left-handed vector-like lepton and the right-handed charged lepton due to the $Z_{2}$ symmetry. Assuming the block matrix $M$ is non-singular, the mass matrix in eq. (2.7) can be block-diagonalized as
[…]
in leading orders of small $3\times 3$ dimensionless perturbation matrices $x\equiv\left(m_{\psi}M^{-1}\right)^{*}$ and $z\equiv m_{E}\left(M^{-1}x^{\dagger}\right)^{T}$, by using the following two rotations:
$\tilde{U}_{L}=\begin{pmatrix}\mathbf{I}_{3}-\frac{1}{2}x\,x^{T}&x\\[5.0pt] -x^{T}&\mathbf{I}_{3}-\frac{1}{2}x^{T}\,x\end{pmatrix},\quad\tilde{V}_{R}=\begin{pmatrix}\mathbf{I}_{3}-\frac{1}{2}z\,z^{T}&z\\[5.0pt] -z^{T}&\mathbf{I}_{3}-\frac{1}{2}z^{T}\,z\end{pmatrix},$ (2.9)
where $\mathbf{I}_{3}$ and $\mathbf{O}_{3}$ are three dimensional identity matrix and zero matrix, respectively. Furthermore, each blocks in eq. (2.8) are further diagonalized by bi-unitary rotations,
$U_{L}^{(e)\dagger}m_{E}V^{(e)}_{R}=\operatorname{diag}(m_{1},m_{2},m_{3}),$ (2.10)
$U_{L}^{(\psi)\dagger}\left(M+\frac{1}{2}x^{\dagger}\,m_{\psi}\right)V^{(\psi)}_{R}=\operatorname{diag}(M_{1},M_{2},M_{3}).$
Here, the notations $m_{i}$ and $M_{i}$, respectively, denote the charged lepton and the vector-like lepton masses. Ultimately, complete diagonalization of the mass matrix in eq. (2.7) is achieved by
$U_{L}^{\dagger}M_{E}V_{R}=\operatorname{diag}(m_{i},M_{i}),$ (2.11)
where $U_{L}=\tilde{U}_{L}\left(U_{L}^{(e)}\oplus U_{L}^{(\psi)}\right)$ and $V_{R}=\tilde{V}_{R}\left(V_{R}^{(e)}\oplus V_{R}^{(\psi)}\right)$.
In the neutral lepton sector, the mass matrix in the bases of $\left(\overline{\nu_{i,L}},\overline{\psi^{0}_{i,L}}\right)$ and $(\nu_{j,R},\psi^{0}_{j,R})^{T}$ is given by
$M_{\nu}=\begin{pmatrix}0&m_{\psi}\\ m_{D}&M\end{pmatrix},$ (2.12)
where the Dirac mass matrix, $m_{D}$, is defined as $m_{D}\equiv\frac{1}{\sqrt{2}}Y_{\nu}v$, mirroring the structure observed in the charged lepton sector. Each block of the neutral lepton mass matrix consists of $3\times 3$ submatrices. The diagonalization procedures similar to those outlined in eqs. (2.8)-(2.11) are applied here, and summarized as follows. The initial block diagonalization
[…]
is facilitated by rotation matrices $\tilde{\mathcal{U}}_{L}$ and $\tilde{\mathcal{V}}_{R}$ with small correction terms $\epsilon=\left(m_{\psi}M^{-1}\right)^{*}$ and $\eta=\left(M^{-1}m_{D}\right)^{T}$, detailed as
$\tilde{\mathcal{U}}_{L}=\begin{pmatrix}\mathbf{I}_{3}-\frac{1}{2}\epsilon\,\epsilon^{T}&\epsilon\\[5.0pt] -\epsilon^{T}&\mathbf{I}_{3}-\frac{1}{2}\epsilon^{T}\epsilon\end{pmatrix},\quad\tilde{\mathcal{V}}_{R}=\begin{pmatrix}\mathbf{I}_{3}-\frac{1}{2}\eta\,\eta^{T}&\eta\\[5.0pt] -\eta^{T}&\mathbf{I}_{3}-\frac{1}{2}\eta^{T}\eta\end{pmatrix}.$ (2.14)
The top-left block in eq. (2.13) contains the light neutrino mass matrix
$m_{\nu}=-m_{\psi}M^{-1}m_{D}=-\frac{1}{2}Y_{\psi}M^{-1}Y_{\nu}(uv).$ (2.15)
Each block in eq. (2.13) is subsequently diagonalized by $3\times 3$ unitary matrices $\mathcal{U}_{L}^{(\nu)}$, $\mathcal{V}^{(\nu)}_{R}$, $\mathcal{U}_{L}^{(\psi)}$, $\mathcal{V}^{(\psi)}_{R}$ that rotate the left- and right-handed neutral leptons, yielding
$\mathcal{U}_{L}^{(\nu)\dagger}m_{\nu}\mathcal{V}^{(\nu)}_{R}=\operatorname{diag}\left(m^{(\nu)}_{1},m^{(\nu)}_{2},m^{(\nu)}_{3}\right),$ (2.16)
$\mathcal{U}_{L}^{(\psi)\dagger}\left[M+\frac{1}{2}m_{D}\left(M^{-1}m_{D}\right)^{T}\right]\mathcal{V}^{(\psi)}_{R}=\operatorname{diag}\left(M^{(0)}_{1},M^{(0)}_{2},M^{(0)}_{3}\right).$ (2.17)
Finally, masses of the light neutrinos and the neutral components of the vector-like leptons are achieved by full diagonalization of the mass matrix
$\mathcal{U}_{L}^{\dagger}M_{\nu}\mathcal{V}_{R}=\operatorname{diag}\left(m^{(\nu)}_{i},M^{(0)}_{i}\right),$ (2.18)
$\mathcal{U}_{L}=\tilde{\mathcal{U}}_{L}\left(\mathcal{U}_{L}^{(\nu)}\oplus\mathcal{U}_{L}^{(\psi)}\right),\quad\mathcal{V}_{R}=\tilde{\mathcal{V}}_{R}\left(\mathcal{V}_{R}^{(\nu)}\oplus\mathcal{V}_{R}^{(\psi)}\right).$
Notice that this approach highlights the mechanism by which the small Dirac neutrino masses arise, featuring a double suppression through the large mass scale $M$ and small Yukawa couplings $Y_{\psi}$.
3 Phenomenological analysis
The previous section detailed the origins of Dirac neutrino mass, explaining its smallness through a vector-like lepton mediated seesaw mechanism. This section focuses on the phenomenological consequences, encompassing concrete realizations of neutrino masses, mixing, leptonic CP violation, and the explanation of dark matter.
Neutrino masses are derived from the mass matrix in eq. (2.15) through the diagonalization procedure outlined in eq. (2.16). Before settling on a basis, it is evident that there are three basis-independent constraints between the neutrino mass matrix elements and the squares of the neutrino masses:
$\operatorname{Tr}(m_{\nu}m_{\nu}^{\dagger})=\sum_{i,j=1}^{3}\left|m_{\nu,ij}\right|^{2}=\sum_{i=1}^{3}\left[m^{(\nu)}_{i}\right]^{2},$ (3.1)
$\det(m_{\nu}m_{\nu}^{\dagger})=\left|\det(m_{\nu})\right|^{2}=\prod_{i=1}^{3}\left[m^{(\nu)}_{i}\right]^{2},$ (3.2)
$\frac{1}{2}\left(\operatorname{Tr}[m_{\nu}m_{\nu}^{\dagger}]^{2}-\operatorname{Tr}[(m_{\nu}m_{\nu}^{\dagger})^{2}]\right)=$ $\frac{1}{2}\sum\varepsilon_{ijk}^{2}\varepsilon_{abc}^{2}\left|m_{\nu,jb}\right|^{2}\left|m_{\nu,kc}\right|^{2}$ (3.3)
$-\sum\varepsilon_{ijk}^{2}\mathrm{Re}\left(2m_{\nu,ik}m_{\nu,jj}m_{\nu,ij}^{*}m_{\nu,jk}^{*}+m_{\nu,ii}m_{\nu,jj}m_{\nu,ij}^{*}m_{\nu,ji}^{*}\right)$
$=\left[m^{(\nu)}_{1}m^{(\nu)}_{2}\right]^{2}+\left[m^{(\nu)}_{1}m^{(\nu)}_{3}\right]^{2}+\left[m^{(\nu)}_{2}m^{(\nu)}_{3}\right]^{2},$
where all summed indices range from 1 to 3. These equations not only remain independent of the basis choice but also provide connections between particular combinations of mass matrix elements on the left and functions of the physical neutrino masses on the right. Specifically, eqs. (3.1),  (3.2) and  (3.3) are the linear, cubic, and quadratic invariants of the neutrino mass squares, respectively. For instance, eq. (3.1) indicates that the sum of the modulus squares of the mass matrix elements should not exceed the sum of the squares of the three light neutrinos’ masses. According to recent global fit results for neutrino oscillation [46, 47, 48, 49], even the largest modulus square of the mass matrix cannot exceed $\mathcal{O}(10^{-3})$ eV2, assuming the lightest neutrino mass is zero.
While this provides a qualitative overview of the small neutrino mass generation, a similar flavor issue as in the SM persists. The relative sizes of these matrix elements are not inherently constrained by the theory; rather, they are restricted by experimental results. All permissible elements can vary widely, leaving the observed hierarchy in lepton masses and the form of their mixing matrix unexplained. This issue could be circumvented by introducing a flavor symmetry, which would shape the forms of mass matrices. For illustrative purposes, one might consider the implementation of models such as the $A_{4}$ flavor symmetry [50, 51], or explore scenarios involving extra-dimensional warped geometries [52]. For instance, under an $A_{4}$ symmetry, the charged lepton mass matrix could be constrained to the following form
$m_{E}=\frac{v}{\sqrt{2}}\begin{pmatrix}y_{e}&y_{\mu}&y_{\tau}\\ y_{e}&\omega y_{\mu}&\omega^{2}y_{\tau}\\ y_{e}&\omega^{2}y_{\mu}&\omega y_{\tau}\end{pmatrix},$ (3.4)
which can be diagonalized, $U_{L}^{(e)\dagger}m_{E}m_{E}^{\dagger}U_{L}^{(e)}=\operatorname{diag}(m_{e}^{2},m_{\mu}^{2},m_{\tau}^{2})$, by the unitary matrix
$U_{L}^{(e)}=\frac{1}{\sqrt{3}}\begin{pmatrix}1&1&1\\ 1&\omega&\omega^{2}\\ 1&\omega^{2}&\omega\end{pmatrix},$ (3.5)
yielding lepton masses $m_{\alpha}=\sqrt{3/2}\left|y_{\alpha}\right|v$ for $\alpha=e,\mu,\tau$. Such symmetry-driven constraints not only provide a theoretical basis for the mass hierarchy and mixing patterns observed experimentally, but also offer a structured pathway for future theoretical and experimental investigations.
Given the neutrino mass matrix, working in the mass basis of the vector-like lepton, it can be constructed as follows:
$Y_{\psi}=-rY_{\nu}^{T},\quad M=\operatorname{diag}(M_{1},M_{2},M_{3}),\quad Y_{\nu}=\begin{pmatrix}y_{11}&0&y_{13}\\ 0&y_{22}&0\\ y_{31}&0&y_{33}\end{pmatrix},$ (3.6)
where $Y_{\psi}$ is supposed to be proportional to the transpose of $Y_{\nu}$ with the incorporation of a small constant factor $r$. The resulting light neutrino mass matrix is given by
$m_{\nu}=\frac{uvr}{2}\begin{pmatrix}\frac{y_{11}^{2}}{M_{1}}+\frac{y_{31}^{2}}{M_{3}}&0&\frac{y_{11}y_{13}}{M_{1}}+\frac{y_{31}y_{33}}{M_{3}}\\ 0&\frac{y_{22}^{2}}{M_{2}}&0\\ \\[0.01pt] \frac{y_{11}y_{13}}{M_{1}}+\frac{y_{31}y_{33}}{M_{3}}&0&\frac{y_{13}^{2}}{M_{1}}+\frac{y_{33}^{2}}{M_{3}}\end{pmatrix}.$ (3.7)
The neutrino masses
$m^{(\nu)}_{1,3}=$ $\frac{uvr}{2\sqrt{2}}\left[\left|\frac{y_{11}^{2}}{M_{1}}+\frac{y_{31}^{2}}{M_{3}}\right|^{2}+\left|\frac{y_{13}^{2}}{M_{1}}+\frac{y_{33}^{2}}{M_{3}}\right|^{2}+2\left|\frac{y_{11}y_{13}}{M_{1}}+\frac{y_{31}y_{33}}{M_{3}}\right|^{2}\mp\sqrt{\triangle}\right]^{1/2},$ (3.8)
$\triangle$=△ = $\left(\left|\frac{y_{13}^{2}}{M_{1}}+\frac{y_{33}^{2}}{M_{3}}\right|^{2}-\left|\frac{y_{11}^{2}}{M_{1}}+\frac{y_{31}^{2}}{M_{3}}\right|^{2}\right)^{2}$
$+4\left|\left(\frac{y_{11}y_{13}}{M_{1}}+\frac{y_{31}y_{33}}{M_{3}}\right)\left(\frac{y_{11}^{*2}}{M_{1}}+\frac{y_{31}^{*2}}{M_{3}}\right)+\left(\frac{y_{13}^{2}}{M_{1}}+\frac{y_{33}^{2}}{M_{3}}\right)\left(\frac{y_{11}^{*}y_{13}^{*}}{M_{1}}+\frac{y_{31}^{*}y_{33}^{*}}{M_{3}}\right)\right|^{2},$
$m^{(\nu)}_{2}=$ $\frac{uvr}{2M_{2}}\left|y_{22}\right|^{2},$
are obtained from diagonalization of the Hermitian matrix $\ m_{\nu}m_{\nu}^{\dagger}$ via a unitary transformation $\mathcal{U}_{L}^{(\nu)}$ acting on left-handed neutrino fields. The unitary rotation matrix is parameterized as
$\mathcal{U}_{L}^{(\nu)}=\begin{pmatrix}\cos\alpha&0&\sin\alpha\ e^{-i\beta}\\ 0&1&0\\ -\sin\alpha\ e^{i\beta}&0&\cos\alpha\end{pmatrix}.$ (3.9)
The angular parameters $\alpha$ and $\beta$ are determined by
$\tan 2\alpha=\frac{2\left|\left(\frac{y_{11}y_{13}}{M_{1}}+\frac{y_{31}y_{33}}{M_{3}}\right)\left(\frac{y_{11}^{*2}}{M_{1}}+\frac{y_{31}^{*2}}{M_{3}}\right)+\left(\frac{y_{13}^{2}}{M_{1}}+\frac{y_{33}^{2}}{M_{3}}\right)\left(\frac{y_{11}^{*}y_{13}^{*}}{M_{1}}+\frac{y_{31}^{*}y_{33}^{*}}{M_{3}}\right)\right|}{\left|\frac{y_{13}^{2}}{M_{1}}+\frac{y_{33}^{2}}{M_{3}}\right|^{2}-\left|\frac{y_{11}^{2}}{M_{1}}+\frac{y_{31}^{2}}{M_{3}}\right|^{2}},$ (3.10)
$\beta=\arg\left[\left(\frac{y_{11}y_{13}}{M_{1}}+\frac{y_{31}y_{33}}{M_{3}}\right)\left(\frac{y_{11}^{*2}}{M_{1}}+\frac{y_{31}^{*2}}{M_{3}}\right)+\left(\frac{y_{13}^{2}}{M_{1}}+\frac{y_{33}^{2}}{M_{3}}\right)\left(\frac{y_{11}^{*}y_{13}^{*}}{M_{1}}+\frac{y_{31}^{*}y_{33}^{*}}{M_{3}}\right)\right].$
Moreover, apart from dependencies on the moduli of the neutrino mass matrix elements, the neutrino masses also exhibit a correlation with the parameter $\alpha$ through the following relationship:
$\triangle=\frac{\left(\left|m_{\nu,33}\right|^{2}-\left|m_{\nu,11}\right|^{2}\right)^{2}}{\cos^{2}2\alpha},$ (3.11)
with no such dependency observed for $\beta$. Figure 2 demonstrates that the neutrino mass-squared differences can be accommodated within the allowed range of $\alpha$ (further discussion on its range is provided below). For these data points, we conducted a numerical scan where the neutrino mass matrix elements were varied randomly within the interval $[0,0.01]$, adhering to the mass-squared difference constraints from the experimentally allowed $3\sigma$ region. As demonstrated, in the range of $\alpha$, theoretical predictions for the solar and atmospheric mass-squared differences are in accordance with experimental constraints within the explored range of $\alpha$.
Figure 2: Scaled neutrino mass-squared differences plotted against the allowed range of $\alpha$. The solar neutrino mass-squared difference (blue points) is multiplied by $10^{5}$ and the atmospheric mass-squared difference (red points) by $10^{3}$. The horizontal cyan and green bands represent their respective current experimentally allowed $3\sigma$ regions for neutrino normal mass ordering.
The lepton mixing matrix, which encapsulates contributions from the unitary rotations of both the charged lepton and neutrino sectors, is expressed as
$U\simeq U_{L}^{(e)\dagger}\mathcal{U}_{L}^{(\nu)}=\left(\begin{array}[]{ccc}\frac{\cos\alpha-\sin\alpha\ e^{i\beta}}{\sqrt{3}}&\frac{1}{\sqrt{3}}&\frac{\cos\alpha+\sin\alpha\ e^{-i\beta}}{\sqrt{3}}\\ \\[0.01pt] \frac{\cos\alpha-\omega\sin\alpha\ e^{i\beta}}{\sqrt{3}}&\frac{\omega^{2}}{\sqrt{3}}&\frac{\omega\cos\alpha+\sin\alpha\ e^{-i\beta}}{\sqrt{3}}\\ \\[0.01pt] \frac{\cos\alpha-\omega^{2}\sin\alpha\ e^{i\beta}}{\sqrt{3}}&\frac{\omega}{\sqrt{3}}&\frac{\omega^{2}\cos\alpha+\sin\alpha\ e^{-i\beta}}{\sqrt{3}}\\ \end{array}\right).$ (3.12)
The magnitudes of the elements in this mixing matrix fit well within their $3\sigma$ experimental ranges as announced by recent studies [46, 47]. This matrix introduces two free parameters, $\alpha$ and $\beta$, while generic form of a lepton mixing matrix is parameterized by three angles and one Dirac CP phase. This reduction in parameters allows for potential predictions or correlations among the neutrino oscillation parameters. By comparing the lepton mixing matrix in eq. (3.12) with the standard parameterization [53], we derived expressions that describe the functional dependencies of the neutrino mixing angles on $\alpha$ and $\beta$, as well as predictions for the Jarlskog invariant [54], which quantifies leptonic CP violation in a rephasing invariant manner:
$\sin^{2}\theta_{12}=\frac{1}{2-\sin 2\alpha\cos\beta},\quad\sin^{2}\theta_{13}=\frac{1}{3}\left(1+\sin 2\alpha\cos\beta\right),$ (3.13)
$\sin^{2}\theta_{23}=\frac{1}{2}\left(1-\frac{\sqrt{3}\sin 2\alpha\sin\beta}{2-\sin 2\alpha\cos\beta}\right),\quad J_{\rm CP}=-\frac{\cos 2\alpha}{6\sqrt{3}}.$
These equations reveal correlations among the mixing angles and the CP violating phase
$\sin^{2}\theta_{12}=\frac{1}{3\cos^{2}\theta_{13}},\quad\tan 2\theta_{23}\cos\delta_{\rm CP}=\frac{\cos 2\theta_{13}}{\sin\theta_{13}\sqrt{2-3\sin^{2}\theta_{13}}}.$ (3.14)
With known values for $\alpha$ and $\beta$, the Dirac CP violating phase is determined by
$\cos\delta_{\rm CP}=\frac{\cos\alpha+\sin\alpha\cos\beta}{\sqrt{1+\sin 2\alpha\cos\beta}}.$ (3.15)
This indicates two special cases where leptonic CP violation is either absent or maximal. For instance, $\alpha=0$ results in $\delta_{\rm CP}=0$, regardless of $\beta$. Conversely, $\beta=0$ sets $\delta_{\rm CP}$ to 00 or $\pi$, independent of $\alpha$. These are the points where no leptonic CP violation appears. The condition $\cot\alpha=-\cos\beta$ implies maximal CP violation with $\delta_{\rm CP}=\pm\frac{\pi}{2}$.
Comparative analysis of theoretical results for the Jarlskog invariant, given in eq. (LABEL:eq:thetasandJcp), restricts $\alpha$ to a range of $-0.60\leq\alpha\leq 0.92$, which aligns with the experimental $3\sigma$ range of $-0.0348\leq J_{\rm CP}\leq 0.0263$ [46, 47]. Additional constraints on $\alpha$ and $\beta$ stem from global fit data regarding mixing angles, CP violating phase, and their correlations. As depicted in figure 3, the theoretical predictions are consistent with experimental results, given the same values and ranges of $\alpha$ and $\beta$. The sharp turn occur in the curves of panels (a) and (b) when $\alpha=\frac{\pi}{4}$ as it is a singular point, c.f. eq.(LABEL:eq:alphabetapara). At this point, the denominator of eq.(3.15) vanishes for $\beta=\pi$. The panel (c) depicts that theoretical result can cover the region which lies above the best fit point.
Figure 3: Plots of the CP-violating phase over $\sin^{2}\theta_{12}$ (shown in (a)), $\sin^{2}\theta_{23}$ (shown in (b)) and $10^{2}\sin^{2}\theta_{13}$ (shown in (c)), for given values of $\alpha$ indicated in the legend, which are taken the same for all plots, and range of $\beta\in[-\pi,-2.25]$. The black stars are the global best-fit points, the blue, orange and green contours are the current $1\sigma$, $2\sigma$ and $3\sigma$ regions, respectively, from the global fit of neutrino oscillation parameters [46, 47].
Figure 4 further demonstrates how the chosen parameter space can successfully accommodate additional correlations among the sine squares of the mixing angles, providing a good match with experimental constraints across the $1-3\sigma$ regions. Panel (a) of the figure illustrates the correlation between $\sin^{2}\theta_{12}$ and $\sin^{2}\theta_{23}$. Notably, the lower boundaries of the curves are constrained by the upper limit of the parameter range for $\beta$, which is established based on other correlated observations within the model. The panel (b) depicts the relationship between $\sin^{2}\theta_{12}$ and $10^{2}\sin^{2}\theta_{13}$. Interestingly, despite varying $\alpha$ and $\beta$ within their respective ranges, all curves overlap due to both $\sin^{2}\theta_{12}$ and $\sin^{2}\theta_{13}$ being dependent on the single parametric combination of $\sin 2\alpha\cos\beta$, as formulated in eq. (LABEL:eq:thetasandJcp). Although taking different values for $\alpha$ and changing $\beta$ in its range, the combination remains the same. This dependency causes all theoretical curves to merge into a single line. Notably, the plots in panel (b) of figure 3 and panel (c) of figure 4 reveal that $\theta_{23}$ predominantly resides within the higher octant. This observation is particularly significant given the ongoing experimental ambiguity in determining the precise octant of the atmospheric mixing angle.
Figure 4: Correlations among the solar, atmospheric, and reactor mixing angles: (a) shows $\sin^{2}\theta_{12}$ versus $\sin^{2}\theta_{23}$, (b) shows $\sin^{2}\theta_{12}$ versus $10^{2}\sin^{2}\theta_{13}$, and (c) shows $10^{2}\sin^{2}\theta_{13}$ versus $\sin^{2}\theta_{23}$. Parameter values (or ranges), plotting styles, and experimentally allowed regions are consistent with those presented in figure 3.
In addition to elucidating the mechanism for neutrino mass and mixing, this model introduces a dark matter candidate crucial for neutrino mass generation. The neutral component of the vector-like lepton, interacting solely through weak interactions, has its lightest right-handed component potentially long-lived. Given that the scalar $\phi$ is heavier than the lightest right-handed component of the lepton, $m_{\phi}>m_{0}$, (where $m_{0}$ is the mass of the lightest $\psi^{0}_{i,R}$), kinematic constraints prevent two-body decay processes, although the neutral component does couple to the neutrino and scalar $\phi$.
Consider the three-body decay process $\psi^{0}_{R}\to\nu hh$ with $\phi$ as a mediator. The decay rate is described by
$\Gamma=\frac{Y^{2}_{\psi}\lambda^{2}}{(16\pi)^{3}}\frac{\cos^{4}\theta u^{2}(m_{0}-2m_{h})}{m_{0}^{2}}\left[\frac{m_{0}^{2}(m_{0}-2m_{h})}{m_{0}^{2}m_{h}-m_{0}m_{\phi}^{2}+m_{h}m_{\phi}^{2}}+\ln\frac{m_{0}m_{\phi}^{2}-m_{h}m_{\phi}^{2}-m_{h}m_{0}^{2}}{(m_{0}-m_{h})(m_{\phi}^{2}-m_{0}^{2})}\right],$ (3.16)
where neutrino mass is neglected. The small couplings $Y_{\psi}$ and $\lambda$, combined with the heavy mass of the decaying particle, result in an extremely low decay rate, extending its potential lifetime beyond the age of the universe. As this particle participates only in the weak interaction, this characteristic classifies it as a Weakly Interacting Massive Particle (WIMP). The relevant annihilation processes of this WIMP are depicted in figure 5.
Figure 5: Tree-level processes contributing to the annihilation of the WIMP DM, where $f$ denotes the SM fermions, and $V$ represents the weak gauge bosons $W$ or $Z$.
The relic abundance of WIMP DM, a well-established concept in cosmology and particle physics, is described by [55]
$\Omega h^{2}=0.1\left(\frac{x_{f}}{20}\right)\left(\frac{g_{*}(m_{0})}{100}\right)^{1/2}\frac{2\times 10^{-26}\text{cm}^{3}\text{s}^{-1}}{\langle\sigma v\rangle},$ (3.17)
where $x_{f}=m_{0}/T_{f}$ is the mass to temperature ratio at the time of freeze-out, $g_{*}$ represents the number of effective entropy degrees of freedom, and $\langle\sigma v\rangle$ is the thermally averaged annihilation cross-section times the relative velocity. Achieving the observed DM relic density $\Omega h^{2}=0.120\pm 0.001$ as reported by Planck [56] requires $\langle\sigma v\rangle=1.72\times 10^{-9}\text{GeV}^{-2}$, which equals to the numerator in eq. (3.17), typical for weak interaction processes.
Lastly, the scalar sector of this model includes a heavy scalar $\phi$, mainly composed of the singlet scalar $\sigma$. As a gauge singlet, $\phi$ lacks direct couplings with SM quarks and gauge bosons, interacting mainly through suppressed mixing with the Higgs. Its primary interactions involve neutrinos, charged leptons, and vector-like leptons [^1] Here, all fields are written in mass basis.,
$\mathcal{L}_{\rm Y}\supset-\cos\theta Y_{\psi,ij}\left[\left(\mathcal{U}_{L}^{(\nu)*}\right)_{ik}\overline{\nu_{kL}}\psi^{0}_{j,R}+\left(U^{(e)*}_{L}\right)_{ik}\overline{e_{kL}}\psi^{-}_{j,R}\right]\phi,$ (3.18)
which is also very weak. Due to such a small interaction strength, $\phi$ is challenging to produce in high energy colliders. Having a significant coupling with the Higgs, $\phi$ decays rapidly into Higgs pairs upon production, making it difficult to detect in high-energy collider environments due to its fleeting presence.
4 Summary and discussion
This paper presents a simple model constructed by extending the SM symmetry with discrete groups $Z_{2}$ and $Z_{3}$, while also expanding the scalar and leptonic sectors of the SM. The model also enlarges the scalar and leptonic part of the SM. The $Z_{2}$ symmetry differentiates between SM fields and newly added fields, while the unbroken $Z_{3}$ symmetry enforces a type of lepton number conservation and ensures the Dirac nature of neutrinos by prohibiting Majorana mass terms. This model exhibits intriguing phenomenological implications, successfully accounting for the normal neutrino mass ordering (see Figure 2), neutrino mixing angles (see Figure 4), leptonic CP violation (see Figure 3), and a WIMP DM candidate.
[…]
As we look to the future, there remain several avenues for further investigation. The inclusion of additional scalars and fermions introduces novel physical phenomena. The model’s Yukawa Lagrangian, featuring flavor-changing couplings, opens possibilities for studying lepton flavor-violating (LFV) processes such as $\mu\to e\gamma$ and/or $\mu\to 3e$. Ongoing and upcoming experiments will impose stringent constraints on these LFV processes, thereby constraining our model’s parameter space. Moreover, the heavy masses of new particles and the complex nature of Yukawa couplings could potentially facilitate Dirac leptogenesis, suggesting an avenue for generating the observed matter-antimatter asymmetry in the universe. Additionally, the collider signatures associated with vector-like leptons, right-handed neutrinos, and additional scalar fields introduced by our model could manifest in observable phenomena at high-energy colliders such as the Large Hadron Collider (LHC) and future facilities. While our current study has employed discrete symmetry groups to maintain simplicity, there is potential to achieve a similar Dirac seesaw model through expansion of the symmetry group or by extending to continuous symmetries, which may provide deeper phenomenological insights. These aspects, while not explored in this paper, are slated for investigation in our future work.
Acknowledgements
The work of Y.R. is supported by the Natural Science Foundation of Xinjiang Uygur Autonomous Region of China under Grant No. 2022D01C52. The work of M.A. is supported by the National Natural Science Foundation of the People’s Republic of China (No. 12303002) and Tianchi talent project of Xinjiang Uygur autonomous region of China.